Atsie (atsie):

Can someone tell me a step by step explanation on what distributive properties are and also give an example?

OpenStudy (skullpatrol):

Do you have any ideas? :-)

OpenStudy (skullpatrol):

The distributive properties state that a multiplier can be distributed over terms. The call it distribut"IVE" because the multiplication CAN be done producing a simplified result. Consider the expression: a • ( b + c ) The multiplier "a" can be distributed over the ( b + c ) producing the simplified result: a • b + a •c

OpenStudy (skullpatrol):

Here is an example: a • ( b + c ) = a • b + a • c 3 • ( 4 + 2 ) = 3 • 4 + 3 • 2 3 • ( 6 ) = 18 also 12 + 6 = 18 18 = 18 so the property is correct

Atsie (atsie):

Finally a response! Thank you! Now whats funny is I am learning this in Algebra class but I'm given a rather different more simple example of a (b + c) = ab + ac and (b+c) a = ba + ca and a (b - c) = ab - ac and (b - c) a = ba - ca. This is about the only example I am given in the book, and there aren't any other examples with more specifics. So I am thrown into the deep end with a test question stating this: "Rewrite the expression using the distributive property. 5 ∙ 97" What are proper steps to solve this?

OpenStudy (anonymous):

im here

OpenStudy (anonymous):

@Atsie im here :)

Atsie (atsie):

ooh, didn't see ya. I'm sure you've read it over right?

OpenStudy (anonymous):

yeah im still

OpenStudy (anonymous):

what is your question?

Atsie (atsie):

Well, firstly have you learned about Distributive Properties yet? I don't want to burden you if you don't know it.

OpenStudy (anonymous):

i have just its kinda rusty in there, but if i read a question im sure ill remember

Atsie (atsie):

Well here is the first page of my chapter, maybe you'll get a better idea. :)

OpenStudy (anonymous):

im slowly getting it

OpenStudy (anonymous):

sorry im taking long

Atsie (atsie):

Its ok. if you need the second page let me know.

OpenStudy (anonymous):

yeah i thinnk i do

Atsie (atsie):

Here you go

OpenStudy (anonymous):

oh i get it!

OpenStudy (anonymous):

ok whats the question?

Atsie (atsie):

Rewrite the expression using the distributive property. 5 ∙ 97

Atsie (atsie):

I don't know if its just me, but everything in the book seems soo simple to the point of ridculous and then I get this for a question and it seems so complicated

OpenStudy (anonymous):

first do 5 x 97

Atsie (atsie):

thats 485

OpenStudy (anonymous):

ok now the thing it 5(485)

OpenStudy (anonymous):

is*

Atsie (atsie):

Right, so what is the equation getting to that?

Atsie (atsie):

I mean, I'm given these options to choose from....

Atsie (atsie):

A. (5 + 90)7 B. 5(100 - 3) C. (10 - 5)7 D. 5(100 - 17)

OpenStudy (anonymous):

is one of the answres 2425?

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

oh! it is A sorry for direct answere

OpenStudy (anonymous):

answer

Atsie (atsie):

Ok thats good to know, but how is it that though? I mean I believe you, but I'm wondering how it makes sense 5 + 90 is 95. Do I do 95 x 7?

Atsie (atsie):

Wait no, that don't even make sense. IF I do that it comes out to 665

OpenStudy (anonymous):

dang now im confused again...

OpenStudy (anonymous):

yeah

Atsie (atsie):

(5 + 90)7 The parentheses mean it leads to a different equation I think. Ugh, I'm lost. You see, this is why the book doesn't seem to be much help to me. I mean its all simple and great in there, but when it comes to actually testing a student on it it just don't make sense!

OpenStudy (anonymous):

look at 2nd page it helped me a bit

OpenStudy (anonymous):

ohhh it B

OpenStudy (anonymous):

becase 5(100 - 3) 100 - 3 is 97 and well its hard to explain

OpenStudy (anonymous):

@Atsie

Atsie (atsie):

I see you I'm trying to re-read the second page and make sense outa those examples. 100-3 does equal 97, that sounds more realistic but I'm unsure of what the 5 in that option has to do with it. Or is that just there to give an idea

OpenStudy (anonymous):

Sorry.. if you don't trust me its ok, But i really think it is B and it is hard to explain but if you dont believe me.. then.. Ok sorry im not too help full

Atsie (atsie):

I believe you, I'm just confused. I sometimes over analyze stuff. I'll go ahead and pick it.

Atsie (atsie):

Ugh, I got nine more of these questions left. Oh joy

OpenStudy (anonymous):

next one i get it now and its easy

OpenStudy (anonymous):

Atsie do you do k12?

Atsie (atsie):

Rewrite the expression using the distributive property. 8 ∙ 203 I'm supposed to be in 10th grade but I got held behind. I've never ever been good at math, I'm usually a failure at it heh

OpenStudy (anonymous):

ok lemme quikly do it can you put the answeres?

Atsie (atsie):

A. 8(20 + 3) B. (8 + 20)3 C. 8(200 + 3) D. (800 + 2)3 Its C isn't it?

OpenStudy (anonymous):

its C sorry again for direct

Atsie (atsie):

Yass, you know what I think you just gave me an idea of what I'm supposed to be doing on this.

OpenStudy (anonymous):

Oh, Lol

Atsie (atsie):

Ok, I'm stuck again I think

OpenStudy (anonymous):

show me the question

Atsie (atsie):

Simplify the expression.(4 - 3m) 8 A. 32 - 24m B. 4 - 24m C. 24m - 32 D. 24m - 4

OpenStudy (anonymous):

im lost..

Atsie (atsie):

Is it A

OpenStudy (anonymous):

Oh yeah it is!

Atsie (atsie):

Cuz 32 - 24 is 8

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

next question?

Atsie (atsie):

It gets more complicated as I go along Simplify the expression 4b - 1 + 5b A. 8b B. -1b + 1 C. -1b - 1 D. 9b - 1

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

sorry but i have to go..

OpenStudy (anonymous):

not trying to avoid but i gtg

OpenStudy (anonymous):

Bye thanks!

Atsie (atsie):

Its ok, I had to go to. Thanks for all your help! I really really appreciate it. I've been stuck for months on end

OpenStudy (skullpatrol):

The numerical examples you are struggling with illustrate how you can use the distributive property to simplify mental arithmetic. They are trying to show that if you want to multiply two numbers together the distributive property allows you to write one of the numbers as a sum or difference and then distibute the other number over the two terms.

OpenStudy (skullpatrol):